Shift Summary for Owl shift, 1/Dec/9

Miguel Barrio (mbarrio@hep.uchicago.edu)
Wed, 1 Dec 1999 07:08:00 -0600

Shift Summary for Owl shift, 1/Dec/99

On Shift: A. Glazov, M. Barrio

Hours of beam available at all intensities: 5.5
Hours of beam available at nominal intensity: 5.5
Hours of E799 data (special runs not included): 5.2
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:
Had to change tapes right at the beginning of the shift.
Got some pipeline errors, had to hard-reboot crate 11, init went fine
after that.

========================================================================
Runs ended this shift:
15100 BEAM_799 - 5.9M evts.

Tapes brought to FCC this shift:
none.

========================================================================
Accesses during shift:
none.

========================================================================
Overview of shift:
E799 datataking, beam was steady, for the most part, but spill quality
wasn't very good: many spills had this big dip (higher intensity) due
to some instabilities in the machine. MCR was informed, they said they
know about this problem and are working on fixing it.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15100 0:01 End of run. Need to change tapes.

15101 0:10 Beginning of run.

0:30 Spill quality has this huge dip (high intensity
"bump") which fires off some live time and
other scalers' alarms.

0:55 Called MCR, told them that "This ain't pretty."
They said they know about the problem and are
working on it.

2:27 Beams goes away, TeV quench or somethin'.

2:42 MCR message: 1hr estimate for beam.

3:54 Beam comes back.

5:49 Beam goes away, LRF4 went down due to an RF
trip (?).

5:52 Beam's back!

262 6:05 Pipeline errors galore. Tried to reset the
pipeline, but crate 11 refuses to respond
hanging up the whole pipe.tcl script.

273 6:15 We do a hard-reset of crate 11 and after that
the init goes fabulously and we're back onto
data taking.

========================================================================
Follow up Notes on Previous Shift
No muon run, you slackers!

========================================================================
Items to Follow up on Next Shift
The scaler errors that will bug you to no end are correlated with the
high intensity dips in the spill structure. Call MCR and bitch and
moan about it... Maybe they'll listen to ya and do something about it.

-- 
Miguel Barrio
Email:  mbarrio@hep.uchicago.edu
Office: HEP 319, 773-702-7484; fax X-1914
Pager:  630-722-6854